# The Learning Approach to Games

### v2

@@ 2 Discrete Games
    $\vec{\varphi}$ are the priors of players,
@@
    $\int_{\hat{\Omega}}\int_{{\cal A}^{i}}\Big{(}\ \sup_{\tilde{\alpha}\in{\cal A}^{i}}{}^{n}J^{i}(\hat{\omega},t,x,\tilde{\alpha})-{}^{n}J^{i}(\hat{\omega},t,x,\alpha)\ \Big{)}{}^{n}\hat{\pi}^{i}(\hat{\omega},t,x)(d\alpha)\hat{\mathbb{P}}(d\hat{\omega})\leq\varepsilon,\quad\forall i\in\mathbb{N}_{0},n\in\mathbb{N}$
  • (iii)
    $\mathbb{P}^{u}\Big{(}\operatorname*{liminf}_{n\to\infty}\sup_{i\in\mathbb{N}_{0}}d^{t,x;i}({}^{0}{\Upsilon}_{\mathfrak{L}}^{t,x;i},{}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x;i})>r\Big{)}\leq\delta$
    where $d^{t,x;i}$ is the metric player $i$ equips the space ${\cal P}({\cal A}^{i})$ under the equivalence $=^{t,x;i}$.
We say $\vec{\varphi}\in\mathscr{M}_{\varphi}^{{\bf N}_{0}}$ is an $(\varepsilon,r,\delta)$-uncertain equilibrium if it is an $(\varepsilon,r,\delta)$-uncertain equilibrium for all $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$.
@@
We remark that condition (iii) aligns with the abstract notion of convergence defined in the Introduction. We have presented it by relying on the priors of players, as this approach is more intuitive and aligns better with the established notations. In essence, we are imposing a further condition (ii) on estimations to achieve a more favorable concept of equilibrium. Later in this section, we will introduce an additional learning parameter and discuss how to incorporate it into this definition.
@@
Prior to discussing the definition, let us revisit the example of chess again. Consider a well trained chess player, and suppose the game is nearing its end. At this late stage, there are potentially many configurations where the subsequent moves are certain. That is, a particular action has an induced probability of $1$, and it remains unchanged as the player continues to learn. We then say that the player is in equilibrium at those particular end game configurations. Similarly, at the very beginning of the game, the player might have a distribution over different openings. Although we will not observe the same opening in each game, for a well-trained player, the distribution over these openings may evolve only over long time scales. We also recognize this as an equilibrium. On the other hand, there may be many other configurations where learning ever continues.
@@
To briefly elaborate on how players solve their own optimization problems, this is partly achieved by repeating past observations with evolving estimations. As player learns, that is as $n$ increases, and recalls past observations, the exploration of other potential scenarios under the current estimations ${}^{n}\hat{\varphi}^{i}$ aims to capture the term $\sup_{\alpha\in\mathcal{A}^{i}}{}^{n}J^{i}(\hat{\omega},t,x,\alpha)$. During this revaluation and exploration process, new strategies may be discovered, or a new assessment of value might lead to changes in ${}^{n}\hat{\pi}^{i}$. Condition (ii) in the definition of uncertain equilibrium implies that the player has explored potential strategies and is capable of generating the best ones under various scenarios of $\hat{\Omega}$ as learning continues.
@@
On the other hand, the values of the player are driven externally. Not only is there no universal notion of what holds higher value, but values often involve multiple, conflicting objectives. These are shaped by needs, interactions, and self-evaluations, as reflected in the remarkable diversity of values across individuals and societies. We would like to emphasize that the equilibrium can be understood specifically for each $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$. With this context in mind, as the player considers a past state that could potentially recur, condition (iii) in the definition of uncertain equilibrium implies that the induced distribution over strategies also recurs as learning progresses.
The choice of $\operatorname*{liminf}$ instead of $\operatorname*{limsup}$ is important. Let us consider two scenarios: (i) Players might remain in a particular equilibrium for nearly all of the time as learning progresses, but they may occasionally explore other actions. Such behavior is observable even in the simplest settings, as we will demonstrate later. We still recognize this as an equilibrium. The use of $\operatorname*{liminf}$ implies that the players are within an $r$-distance of this particular equilibrium infinitely often. (ii) There might be two or more equilibria between which players switch. We recognize each one as an equilibrium, as players will find themselves in each particular equilibrium infinitely often. If we were to require $\operatorname*{limsup}$ instead, it would imply that players are stuck in a particular equilibrium indefinitely, which is quite hard to expect in reality.
@@
Let us briefly explain the role of $(\varepsilon,r,\delta)$, which we took uniform over players for simplicity. Firstly, $\varepsilon$ characterizes how effectively players can solve their targeted optimization problem. Next, $r$ measures how closely the distribution of strategies are approaches the equilibrium infinitely often. Lastly, $\delta$ reflects the likelihood that an equilibrium will be observed. Recall that $\mathbb{P}^{u}$ is associated with the universe in which the players exist; this may depend on the real underlying dynamics, random choices generated by each player etc.
Let us emphasize again that the concept of equilibrium is characterized by the recurrence of the induced strategy distributions of players. In fact, we remark that ${}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x;i}$ can be viewed as a function ×$\mathbb{N}\times\Omega^{u}\to{\cal P}({\cal A}^{i})$, thus a discrete time stochastic process taking values in the space of probability distributions over ${\cal A}^{i}$. Therefore, the uncertain equilibrium is precisely a recurrent point if $r,\delta=0$. The primary interest lies in understanding the evolution of ${}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x;i}$ under a specific learning algorithm, which characterizes the observed behaviors. This offers an alternative perspective for relaxed controls. Indeed, we aim to consider random strategies because the values are inherently uncertain. This is again familiar as better comprehension of the values allows for the construction of more certain strategies.
@@
Now, we will clarify the similarities and differences with the concept of correlated equilibrium. To do so, let us look at the optimality condition in the definition of uncertain equilibrium 3 as follows, simplifying notation by omitting $n$ and $(t,x)$:
$\int_{\hat{\Omega}}\int_{{\cal A}^{i}}J^{i}(\hat{\omega},\alpha)\hat{\pi}^{i}(\hat{\omega})(d\alpha)\hat{\mathbb{P}}(d\hat{\omega})$
$=\int_{\hat{\Omega}}\int_{{\cal A}^{i}}\int_{\vec{\cal A}}J^{i}(\hat{\omega},\vec{\alpha})\hat{\Gamma}^{i}(\alpha;d\vec{\alpha})\hat{\pi}^{i}(\hat{\omega})(d\alpha)\hat{\mathbb{P}}(d\hat{\omega})$
and one notices that
$\int_{\hat{\Omega}}\int_{{\cal A}^{i}}\hat{\Gamma}^{i}(\alpha;d\vec{\alpha})\hat{\pi}^{i}(\hat{\omega})(d\alpha)\hat{\mathbb{P}}(d\hat{\omega})\in{\cal P}(\vec{\cal A})$ (2.12)
Note that the $i$-th marginal of (2.12) is ${\Upsilon}^{i}$, the induced distribution of player $i$. To clarify the connections, consider any $\rho\in{\cal P}(\vec{\cal A})$. By disintegrating $\rho$ onto the $i$-th component as $\rho(d\vec{\alpha})=\rho^{-i}(d\vec{\alpha}|\alpha^{i})\rho^{i}(d\alpha^{i})$, we have that $\rho^{-i}$ corresponds to $\hat{\Gamma}^{i}$ [^7] One can extend $\rho^{-i}$ to the full $\vec{\cal A}$ by the Dirac distribution on $\alpha^{i}$ for the $i$-th marginal, which is also the case for $\hat{\Gamma}^{i}$. and $\rho^{i}$ corresponds to $\hat{\pi}^{i}(\hat{\omega})(d\alpha)\hat{\mathbb{P}}(\hat{\omega})$. Roughly speaking, equilibrium conditions are (with simplified notations),
@@
Uncertain Equilibrium: $\int_{\hat{\Omega}}\sup_{\tilde{\alpha}^{i}}\int_{\vec{\cal A}}J^{i}(\hat{\omega},\tilde{\alpha}^{i},\vec{\alpha}^{-i})\rho^{-i}(d\vec{\alpha}|\tilde{\alpha}^{i})\hat{\mathbb{P}}(d\hat{\omega})$
@@
where the supremum over $\tilde{\alpha}^{i}$, for the Nash-type equilibrium, depends on $\vec{\alpha}^{-i}$; for the correlated equilibrium, it depends on $\alpha^{i}$; for the uncertain equilibrium, it depends on $\hat{\omega}$; and for the coarse correlated equilibrium, it is independent of the controls. Note that we have changed $\hat{\pi}$ to $\sup_{\tilde{\alpha}^{i}}$ due to the optimality condition, and $\hat{\mathbb{P}}$ plays the same role as $\rho^{i}$. However, the key difference is that the correlated equilibrium, similar to the Nash equilibrium, considers deviations without affecting other players, whereas in uncertain equilibrium, changing the strategy would affect others.
Recall that if regret is defined similarly to the correlated equilibrium and is assumed to be sublinear, then the observed time-averaged joint distribution over actions converges to a correlated equilibrium. Considering the estimated distributions over strategies in (2.12), one might expect that all players will eventually induce the same distribution in symmetric situations. However, in general, there is no reason to assume that a single distribution will characterize every player’s considerations.
The concept of Nash equilibrium focuses solely on controls, according to their associated scalar values, inherently excluding the intrinsic structure of a player. For example, in situations where a central planner announces policies for individual agents, such as environmental regulations, traffic management, public health initiatives, with the knowledge that every individual will act according to their own assessment of value, Nash equilibrium is definitely the right consideration. The central planner needs to model agent’s individual values to construct stable policies. In such cases, it is not meaningful to model each and every agent in detail of their learning algorithms. We also remark that, the Nash equilibrium requires players to have exact knowledge of the strategies of other players. In a pure equilibrium sense, this instability is significant enough that an equilibrium does not exist even for simplest games like rock-paper-scissors. To address this, one typically adopts relaxed controls, which is essential but still omits details about the underlying player.
Let us highlight the importance of incorporating additional learnable parameters into our framework. This could include encoding raw observations, establishing communication protocols, and many other spaces of estimations to design more sophisticated player. One such crucial parameter is the players’ expectations regarding the best achievable outcomes, which define a notion of regret for the player and alter the characteristics of exploration, as demonstrated in a simplified form in Section 2.1. Consider functions of the form
@@
${}^{n}\kappa^{i}(t,x):=\hat{\mathbb{P}}\bigg{(}\int_{{\cal A}^{i}}{}^{n}J^{i}(\hat{\omega},t,x,\alpha){}^{n}\hat{\pi}^{i}(\hat{\omega},t,x)(d\alpha)>{}^{n}\hat{B}^{i}(t,x)\bigg{)}$
To relate this quantity to familiar concepts with which we all can relate, we say that at the state $(t,x)\in\mathbb{T}\times\Omega$, the player $i$ is currently desperate if ${}^{n}\kappa^{i}(t,x)=0$, and euphoric if ${}^{n}\kappa^{i}(t,x)=1$. If the player is desperate, as the learning progresses, either ${}^{n}\hat{B}^{i}$ will decrease, leading the player, in some sense, to accept the situation; or ${}^{n}J^{i}$ will assign higher values to underexplored strategies, encouraging the player to explore them. One can describe the player’s current situation in more detail by using verbal subcategories like
@@
which can be set as partitions of $\kappa$ values. Beyond better characterization of a player, we can incorporate this into the definition of equilibrium 3 by requiring
  • (ii’)
    ${}^{n}\kappa^{i}(t,x)>\kappa,\quad\forall i\in\mathbb{N}_{0},n\in\mathbb{N}$
and denoting it as $(\kappa,\varepsilon,r,\delta)$-uncertain equilibrium. That is, we are now searching for an equilibrium where each player is, let’s say, at least confident.
As well as the design of the player, choices for the concept of equilibrium are also diverse. For example, to recover the concept of Nash equilibrium, we can further assume
  • (ii”)
    $\hbox{\rm supp$\,$}\Big{(}{}^{n}\hat{\Gamma}_{t,x,\alpha}^{i}\Big{)}\subset\hbox{\rm supp$\,$}\Big{(}{}^{n}{\Upsilon}^{t,x,1}\times{}^{n}{\Upsilon}^{t,x,2}\times\cdots\Big{)},\quad\forall i\in\mathbb{N}_{0},n\in\mathbb{N},\alpha\in\hbox{\rm supp$\,$}{}^{n}{\Upsilon}^{t,x,i}$
where we extended $\hat{\Gamma}$ to depend on $x$ naturally, and $\times$ denotes the product measure. That is, players’ estimates are indeed within the support of the induced distributions of all the players. If every estimate is predetermined and no randomness is involved, then (ii”) means that players are able to observe each other’s future strategies. Thus, together with (ii) in Definition 3, we recovered the standard Nash equilibrium.
Next, time-consistency, or Dynamic Programming Principle, can be naturally introduced for any learning parameter and can likewise be required at equilibrium. The most important and familiar one is for the value function, and it also relates $\hat{\phi}$ to the concept of the standard value function in the literature. We say that the estimate $(\hat{T},\hat{p},\hat{\Gamma},\hat{F},\hat{\phi},\hat{\pi})$ yields a time consistent value almost surely, if for any $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$ and $0\leq T_{0}\leq\hat{T}(t,x)$, it holds
$\int_{{\cal A}^{i}}J(T_{0};\hat{\omega},t,x,\alpha)\hat{\pi}(\hat{\omega},t,x)(d\alpha)=\int_{{\cal A}^{i}}J(\hat{\omega},t,x,\alpha)\hat{\pi}(\hat{\omega},t,x)(d\alpha)\quad d\hat{\mathbb{P}}-a.s.$ (2.14)
where $J(T_{0};\hat{\omega},t,x,\alpha)$ is defined exactly as in (2.6), where $\hat{T}$ is replaced by $T_{0}$. Notice that, setting $T_{0}=t$ and in case $\hat{\pi}$ yields the optimal control per $\hat{\omega}\in\hat{\Omega}$, (2.14) becomes
$\hat{\phi}(t,x)=\sup_{\alpha}J(t,x,\alpha)=\sup_{\alpha}\int_{\vec{\cal A}}\mathbb{E}^{t,x,\vec{\alpha}}\Big{[}\hat{\phi}(t+\hat{T},X_{t+\hat{T}})+\sum_{s=t}^{t+\hat{T}-1}\hat{F}(s,X_{s},\vec{\alpha}(s,X_{s}))\Big{]}\hat{\Gamma}_{t}(\alpha;d\vec{\alpha})$
@@
To illustrate how time-consistency can be required of an estimation, we focus on $\hat{\pi}$; similar conditions can be formulated for $\hat{T}$, $\hat{p}$, and $\hat{\Gamma}$ as well. We say that the estimate $(\hat{T},\hat{p},\hat{\Gamma},\hat{F},\hat{\phi},\hat{\pi})$ yields a time-consistent distribution over controls almost surely at $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$, if for any $t\leq T_{0}\leq t+\hat{T}(t,x)-1$, and $x_{T_{0}}\in\mathbb{S}_{T_{0}}$, $\hat{\pi}(\hat{\omega},t,x)$ induces the same distribution as $\hat{\pi}(\hat{\omega},T_{0},x_{T_{0}})$ on the space ${\cal A}^{\hat{T}(t,x);T_{0},x_{T_{0}},i}$. Here, the quotient space is defined similarly, with the relation terminating at $t+\hat{T}(t,x)-1$ instead of $t+\hat{T}(T_{0},x_{T_{0}})-1$.

### v4

@@ 3 Discrete Games
    $\hskip-20.00003pt\operatorname*{limsup}_{n\to\infty}\sup_{i\in\mathbb{N}_{0}}\int_{\hat{\Omega}}\Big(\ \sup_{\tilde{\alpha}\in{\cal A}^{i}}{}^{n}J^{i}(\hat{\omega},t,x,\tilde{\alpha})-{}^{n}J^{i}(\cdot,{}^{n}\mathfrak{L}_{\pi}^{i})(\hat{\omega},t,x)\ \Big)\hat{\mathbb{P}}(d\hat{\omega})\leq\varepsilon,$
@@
    $\mathbb{P}^{u}\Big(\operatorname*{liminf}_{n\to\infty}\sup_{i\in\mathbb{N}_{0}}d^{i}({}^{*}{\Upsilon},{}^{n}{\Upsilon}^{i})>r\Big)\leq\delta$
@@
Note that condition (iii) aligns with the recurrence defined in Definition 4. We further impose condition (ii) on estimations to obtain a more favorable notion of equilibrium. Later in this section, we will introduce an additional learning parameter and discuss how to incorporate it into this definition.
@@
Let us revisit the example of chess. Consider a well-trained chess player, and suppose the game is nearing its end. At this late stage, there are often configurations in which subsequent moves are certain. That is, a particular action has an induced probability of one, and remains unchanged as the player continues to learn. Similarly, at the opening of the game, the player might have a distribution over different openings. Although we will not observe the same opening in each game, for a well-trained player this distribution may evolve only over long time scales. On the other hand, there may be many configurations where learning continues indefinitely. We remark that $d$ is an arbitrary metric on functions of the form $\{\hat{\Omega}^{i}\times\mathbb{T}\times\Omega\to\mathbb{A}^{i}\}$, which can be designed to reflect these considerations.
Notice that we require players to approach a particular behavior independent of their previous history. Such independence implicitly requires that players explore sufficiently diverse behaviors to be able to reach this equilibrium. Furthermore, since players may be exploitable, they often change their behaviors. However, an equilibrium is one that recurs infinitely often.
@@
To briefly elaborate on how players might solve their own optimization problems, they may do so by revisiting past observations with evolving estimations. As the player accumulates observations (that is, as $n$ increases) and recalls past observations, the exploration of potential scenarios under the current estimations $^{n}\hat{\varphi}^{i}$ aims to capture the term $\sup_{\alpha\in\mathcal{A}^{i}}{}^{n}J^{i}(\hat{\omega},t,x,\alpha)$. During this revaluation and exploration process, new strategies may be discovered, or an updated assessment of value might lead to changes in $^{n}\hat{\pi}^{i}$. Condition (ii) in the definition of uncertain equilibrium implies that players have explored potential strategies and are capable of generating the best ones under various scenarios of $\hat{\Omega}$ as learning continues.
@@
On the other hand, the values of the player are driven externally. Not only is there no universal notion of what holds higher value, but values often involve multiple, conflicting objectives. These are shaped by needs, interactions, and self-evaluations, as reflected in the remarkable diversity of values across individuals and societies. In settings that we design, scalar values are again externally assigned to observations, and the value function is a representation of these assignments.
The choice of $\operatorname*{liminf}$ instead of $\operatorname*{limsup}$ is important. As mentioned in the Introduction and in Section 2, we interpret convergence of behavior as cooperation between players. By using $\operatorname*{liminf}$, we require that a particular behavior remains favorable and is used infinitely often, though not necessarily always.
@@
Let us briefly explain the role of $(\varepsilon,r,\delta)$, which we take to be uniform over players for simplicity. Firstly, $\varepsilon$ characterizes how effectively players can solve their respective optimization problem. Next, $r$ measures how closely the distribution of strategies approaches the equilibrium infinitely often. Lastly, $\delta$ reflects the likelihood that an equilibrium will be observed. Recall that $\mathbb{P}^{u}$ is associated with the universe in which the players exist; this may depend on the real underlying dynamics, and random choices generated by each player.
We remark that the planned behavior $^{n}{\Upsilon}^{i}(\cdot,t,x)$ in (3.12) is a function of the form $\mathbb{N}\times\Omega^{u}\to\mathbb{A}^{t,x;i}$, thus a discrete-time stochastic process taking values in the space of actions available at $(t,x)$. From this perspective, an uncertain equilibrium can be viewed as a recurrent point of this process. The primary interest lies in understanding the evolution of this process under a specific design of player.
@@
We now clarify the similarities to and differences from the concept of correlated equilibrium. To do so, we examine the optimality condition in Definition 7 of uncertain equilibrium (simplifying notation by omitting n n) as follows:
[…]
and since $\hat{\Gamma}^{i}\in{\cal P}(\vec{\cal A})$ (see (3.4)),
$\int_{\hat{\Omega}}\hat{\Gamma}^{i}(\hat{\pi}^{i}(\hat{\omega});d\vec{\alpha})\hat{\mathbb{P}}(d\hat{\omega})\in{\cal P}(\vec{\cal A})$ (3.13)
[…]
@@
[…]
@@
The supremum over $\tilde{\alpha}^{i}$ has different dependence in each equilibrium concept. In the Nash-type equilibrium [^5] We refer to this as a Nash-type equilibrium due to its structure; however, as it is potentially impossible to satisfy, such a concept does not, to our knowledge, appear in the literature., it depends on $\vec{\alpha}^{-i}$; in the correlated equilibrium, it depends on $\alpha^{i}$; in the uncertain equilibrium, it depends on $\hat{\omega}$; and in the coarse correlated equilibrium, it is independent of the controls. We replace $\hat{\pi}$ with $\sup_{\tilde{\alpha}^{i}}$ to reflect the optimality condition, and $\hat{\mathbb{P}}$ plays a role analogous to $\rho^{i}$. However, the key difference is that the correlated equilibrium, similar to the Nash equilibrium, considers deviations that do not affect other players, whereas in uncertain equilibrium, changing one’s strategy influences others via learned estimations.
Recall that if regret, defined analogously to the correlated equilibrium, is sublinear, then the time-averaged empirical distribution of actions converges to a correlated equilibrium. Considering the estimated distributions over strategies in (3.13), one might expect all players to eventually induce the same distribution in symmetric situations. However, in general, there is no reason to expect that a single distribution characterizes every player’s considerations. Moreover, a similar criticism applies to regret definitions, since changing prior actions would typically influence the future strategies of others.
The concept of Nash equilibrium focuses solely on controls, according to their associated scalar values, inherently excluding the intrinsic structure of a player. For example, in situations where a central planner announces policies for individual agents, such as environmental regulations, traffic management, public health initiatives, with the knowledge that every individual will act according to their own assessment of value, the Nash equilibrium is the appropriate framework. The central planner needs to model agents’ individual values to construct stable policies. In such cases, it is not meaningful to model each and every player in detail through their learning algorithms. We also remark that the Nash equilibrium requires players to have exact knowledge of the strategies of others. In the pure equilibrium sense, this instability is significant enough that an equilibrium may fail to exist even in the simplest games. To address this, one typically adopts relaxed controls, which is essential, though not motivated by player design.
Let us highlight the importance of incorporating additional learnable parameters into our framework. This could include encoding raw observations, establishing communication protocols, and many other spaces of estimations to design more sophisticated player. One such simple yet interesting parameter is a player’s expectation regarding the best achievable outcome, which defines a notion of regret for the player and alters the characteristics of exploration, as demonstrated in simplified form in Section 3.1. Consider functions of the form
@@
$^{n}\kappa^{i}(t,x):=\hat{\mathbb{P}}\bigg({}^{n}J^{i}(\cdot,{}^{n}\hat{\pi}^{i})(\hat{\omega},t,x)>{}^{n}\hat{B}^{i}(t,x)\bigg)$
To relate this quantity to familiar concepts with which we all can relate, we say that at the state $(t,x)\in\mathbb{T}\times\Omega$, the player $i$ is currently desperate if $^{n}\kappa^{i}(t,x)=0$, and euphoric if $^{n}\kappa^{i}(t,x)=1$. If the player is desperate, as the learning progresses, either $^{n}\hat{B}^{i}$ will decrease, leading the player in some sense to accept the situation, or $^{n}J^{i}$ will assign higher values to underexplored strategies, encouraging the player to explore them. One can further describe the player’s current situation using verbal subcategories like
$\small\text{Desperate}-\text{Discouraged}-\text{Doubtful}-\text{Determined}-\text{Confident}-\text{Optimistic}-\text{Euphoric}$
which can be interpreted as partitions of $\kappa$-values. Beyond providing a richer characterization of a player, this notion can be incorporated into Definition 7 of equilibrium by requiring
  • (ii’)
    $\operatorname*{limsup}_{n\to\infty}\sup_{i\in\mathbb{N}_{0}}{}^{n}\kappa^{i}(t,x)>\kappa,$
and denoting it as $(\kappa,\varepsilon,r,\delta)$-uncertain equilibrium. In other words, we now search for player designs that further achieve confidence.
In addition to the design of the player, the choice of equilibrium is also diverse. For example, to recover the concept of Nash equilibrium, we can assume
  • (ii”)
    for all $i\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$ large enough,
    $^{n}\hat{\Gamma}^{i}(t,x)=\text{Law}({}^{n}\hat{\pi}^{1}(\cdot,t,x))\times\text{Law}({}^{n}\hat{\pi}^{2}(\cdot,t,x))\times\cdots$
where we extend $\hat{\Gamma}$ to depend naturally on $x$, while removing dependence on $\alpha$. That is, players’ estimates yield the strategies of others, effectively meaning that players are able to observe each other’s future strategies. If every estimate is predetermined as part of the setting and no randomness is involved, then, together with condition (ii) in Definition 7, we recover that the players’ policies $\hat{\pi}^{i}$ form a Nash equilibrium.
@@
$\int_{{\cal A}^{i}}J(T_{0};\hat{\omega},t,x,\alpha)\hat{\pi}(\hat{\omega},t,x)(d\alpha)=\int_{{\cal A}^{i}}J(\hat{\omega},t,x,\alpha)\hat{\pi}(\hat{\omega},t,x)(d\alpha)$ (3.15)
where $J(T_{0};\hat{\omega},t,x,\alpha)$ is defined exactly as in (3.6), but $\hat{T}$ is replaced by $T_{0}$. Notice that when $T_{0}=t$ and $\hat{\pi}$ yields the optimal control for each $\hat{\omega}\in\hat{\Omega}$, (3.15) becomes
$\hat{\phi}(t,x)=\sup_{\alpha}J(t,x,\alpha)$
$=\sup_{\alpha}\int_{\vec{\cal A}}\mathbb{E}^{t,x,\vec{\alpha}}\Big[\hat{\phi}(t+\hat{T},X_{t+\hat{T}})+\sum_{s=t}^{t+\hat{T}-1}\hat{F}(s,X_{s},\vec{\alpha}(s,X_{s}))\Big]\hat{\Gamma}_{t}(\alpha;d\vec{\alpha})$
@@
[…]